OpenStudy (anonymous):

The rectangular solid below has a volume of 5,600 cubic inches, a width of 20 inches, and a height of 8 inches. What is the length (x) of the rectangular solid? A 20 inches B 35 inches C 200 inches D 280 inches

OpenStudy (mathstudent55):

Since the volume of a rectangular solid is \(V = LWH\). then solving for length, you get \(V = LWH\) \(\dfrac{V}{WH} = L\), or \( L = \dfrac{V}{WH} \) Just plug in your values for V, W, and H, and calculate L.

OpenStudy (anonymous):

I knew that

OpenStudy (mathstudent55):

Great. Now you just need to do it.

OpenStudy (anonymous):

Simple algebra here. 5600= (20)(8)(x) So multiple the 20 by 8, and you get 160. 5600/160=x=35